In a mixture of raisins and dates, the ratio by weight of raisins to dates is 7 to 3. How many pounds of raisins will there be in 7 pounds of this mixture?
Answer by stanbon(75887)

In a mixture of raisins and dates, the ratio by weight of raisins to dates is 7 to 3. How many pounds of raisins will there be in 7 pounds of this mixture?
-------------------------------------
7 to 3 is the same as 7x to 3x where 7x is the weight of raisins:
--------
Equation:
7x + 3x = 7 lbs
10x = 7
x = 0.7
-------
lbs of raisins = 7x = 7*0.7 = 4.9 lbs
=============================
Chers,
Stan H.
==================